# Vec-LUT: Vector Table Lookup for Parallel Ultra-Low-Bit LLM Inference on Edge Devices

### v1

@@ 6.3. Combination with Existing Methods
Since Vec-LUT and existing scalar LUT-based methods are optimized for different input lengths (shown in Fig. 12), it is feasible to combine them to build a universal solution for more diverse inference workloads. For example, when processing a single user request, it is optimal to use Vec-LUT during prefilling (parallel), and switch to scalar LUT-based kernels during decoding (usually not parallel), to minimize the end-to-end inference latency.
@@ 7.1. Ultra-Low-Bit LLM Quantization
Large language models (LLMs) have demonstrated remarkable capabilities but require substantial resources for deployment. To mitigate this, low-bit quantization has emerged as a crucial technique to reduce model size while maintaining model accuracy. For example, LLM.int8() (dettmers2022gpt3) and SmoothQuant (xiao2023smoothquant) adopt 8-bit quantization for both weights and activations to save memory and computation. GPTQ (frantar2022gptq) and AWQ (lin2023awq) further compress the LLM weights to 4 and 3-bit with comparable accuracy to floating point baselines, while keeping activations at FP16 due to random outliers. Besides the post-training quantization (PTQ) methods above, quantization-aware training (QAT) methods (e.g., BitDistiller (du2024bitdistiller), EfficientQAT (chen2025efficientqat), and ParetoQ (liu2025paretoq)) further push the bit width boundary to 2-bit and below. Moreover, Microsoft trains the native ternary (1.58-bit) LLM BitNet from scratch (wang2023bitnet; ma2024158bitllm; ma2025bitnet), which is followed by the community with many emerging ternary LLMs (1bitllm2024bitnet; hf1bitllm2024llama3; Falcon3; kaushal2024spectra). However, the mixed-precision challenge of low-bit, especially ternary LLMs, remains unsolved and significantly hinders their practical deployment on current hardware.
@@ 7.2. LUT-based mpGeMM Optimization
To efficiently support low-bit LLM inference without native hardware support of mpGeMM, LUT-based methods are proposed in replacement of conventional MAD-based methods (more in §2.2). For example, LUT-GEMM (park2022lut) first proposes LUT-based mpGeMM for low-bit LLMs on GPU. T-MAC (wei2024tmac) utilizes the table lookup instructions on CPU to accelerate LUT-based mpGeMM, and even outperforms GPU/NPU-based solutions in single-batch decoding. bitnet.cpp (wang2025bitnet), Microsoft’s official inference framework for ternary LLMs, extends T-MAC’s LUT-based kernel design with ternary-specific optimizations and achieves a remarkable inference speed on CPU. Architectural works (e.g., LUT Tensor Core (mo2024lut), TENET (huang2025tenet), and T-SAR (oh2025tsar)) further accelerate LUT-based mpGeMM by co-designing the LUT algorithm and hardware architecture, demonstrating the superior efficiency of LUT-based solutions. However, all of the existing methods adopt the scalar LUT paradigm, which is inefficient for parallel inference, as discussed in §2.2.2.

### v2

@@ 6.3. Combination with Existing Methods
Vec-LUT is primarily optimized for tasks with long inputs and short outputs (e.g., summarization) or parallel generation. For more diverse workloads, Vec-LUT can be combined with scalar LUT-based methods, switching between them based on the input token length to achieve end-to-end speedup (trade-off shown in Fig. 13). For example, when processing a single user request, it is optimal to use Vec-LUT during prefilling (parallel), and switch to scalar LUT-based kernels during decoding (usually not parallel), to minimize the end-to-end inference latency. Switching between vector and scalar LUT requires weight re-layout due to their different tiling strategies, incurring extra latency if converted at runtime or extra memory if pre-converted.
@@ 7.1. Ultra-Low-Bit LLM Quantization
Large language models (LLMs) have demonstrated remarkable capabilities but require substantial resources for deployment. To mitigate this, low-bit quantization has emerged as a crucial technique to reduce model size while maintaining model accuracy. For example, LLM.int8() (Dettmers et al., 2022) and SmoothQuant (Xiao et al., 2023) adopt 8-bit quantization for both weights and activations to save memory and computation. GPTQ (Frantar et al., 2023) and AWQ (Lin et al., 2024) further compress the LLM weights to 4 and 3-bit with comparable accuracy to floating point baselines, while keeping activations at FP16 due to random outliers. Besides the post-training quantization (PTQ) methods above, quantization-aware training (QAT) methods (e.g., BitDistiller (Du et al., 2024), EfficientQAT (Chen et al., 2025a), and ParetoQ (Liu et al., 2025)) further push the bit width boundary to 2-bit and below. Moreover, Microsoft trains the native ternary (1.58-bit) LLM BitNet from scratch (Wang et al., 2023; Ma et al., 2024, 2025), which is followed by the community with many emerging ternary LLMs (1bitLLM, 2024; Mekkouri et al., 2024; Falcon-LLM Team, 2024; Kaushal et al., 2025). However, the mixed-precision challenge of low-bit, especially ternary LLMs, remains unsolved and significantly hinders their practical deployment on current hardware.
@@ 7.2. LUT-based mpGeMM Optimization
To efficiently support low-bit LLM inference without native hardware support of mpGeMM, LUT-based methods are proposed in replacement of conventional MAD-based methods (more in §2.2). For example, LUT-GEMM (Park et al., 2024) first proposes LUT-based mpGeMM for low-bit LLMs on GPU. T-MAC (Wei et al., 2025) utilizes the table lookup instructions on CPU to accelerate LUT-based mpGeMM, and even outperforms GPU/NPU-based solutions in single-batch decoding. bitnet.cpp (Wang et al., 2025), Microsoft’s official inference framework for ternary LLMs, extends T-MAC’s LUT-based kernel design with ternary-specific optimizations and achieves a remarkable inference speed on CPU. Architectural works (e.g., LUT Tensor Core (Mo et al., 2025), TENET (Huang et al., 2025), and T-SAR (Oh et al., 2025)) further accelerate LUT-based mpGeMM by co-designing the LUT algorithm and hardware architecture, demonstrating the superior efficiency of LUT-based solutions. However, all of the existing methods adopt the scalar LUT paradigm, which is inefficient for parallel inference, as discussed in §2.2.2.
@@ 8. Conclusion
Acknowledgements.
This work is supported by NSFC (No.62402280), CPSF (No.2024M761683), Tsinghua University (AIR)-AsiaInfo Technologies (China), Inc. Joint Research Center for 6G Network and Intelligent Computing, and Xiongan AI Institute.
Appendix A Artifact Appendix
A.1. Abstract
The source code for this work is vlut.cpp, available at https://github.com/OpenBitSys/vlut.cpp. Pre-converted GGUF models are hosted on HuggingFace at https://huggingface.co/collections/XXXXyu/vlutcpp. Detailed evaluation instructions (Evaluation.md) and all scripts are in the evaluation/ directory of the repository.
The instructions below focus on a minimal reproduction of vlut.cpp on a single device with one pre-converted model (<<1 hour). Extending to other models and devices is straightforward (same scripts, different MODEL_DIR and DEVICE_NAME). A full comparison against all baselines requires additional setup described in Evaluation.md (>>1 day per device).
A.2. Artifact check-list (meta-information)
  • Algorithm: vector LUT-based mpGeMM
  • Compilation: CMake $\geq$3.14, C++17 (GCC $\geq$9 / Clang $\geq$10)
  • Model: pre-converted GGUF models, available at https://huggingface.co/collections/XXXXyu/vlutcpp
  • Run-time environment: Ubuntu 20.04+ / WSL2 / Android
  • Hardware: x86_64 AVX2 or ARMv8 NEON/SVE CPU; $\geq$4 GB RAM (inference only) or $\geq$24 GB RAM (manual model conversion)
  • Execution: exclusive access to performance cores
  • Metrics: kernel throughput (runs/s), token throughput (tokens/s)
  • Output: CSV summaries, PDF figures via Python plotting
  • Experiments: GeMM benchmark, end-to-end prefilling, end-to-end batched decoding
  • How much disk space required (approximately)?: $\geq$16 GB for vlut.cpp models; $\geq$64 GB for full comparison with baselines
  • How much time is needed to prepare workflow (approximately)?: $\sim$30 min (build + model download)
  • How much time is needed to complete experiments (approximately)?: <<1 hour for minimal reproduction; >>1 day per device for full comparison
  • Publicly available?: Yes
  • Code licenses (if publicly available)?: MIT
A.3. Description
A.3.1. How to access
A.3.2. Hardware dependencies
  • Minimum: x86_64 CPU with AVX2, or ARMv8 CPU with NEON/SVE, $\geq$4 GB RAM
  • Paper devices: Intel Core i7-13700K (PC), AMD Ryzen 7 5800H (laptop), RK3588/Cortex-A76 (Orange Pi 5 Plus SBC), Snapdragon 8 Elite (Xiaomi 15), ARM Neoverse-V1 (AWS Graviton 3)
A.3.3. Software dependencies
  • OS: Linux (Ubuntu 20.04+), Android (Termux), or WSL2
  • Build: CMake $\geq$3.14, C++17 (GCC $\geq$9 or Clang $\geq$10)
  • Python $\geq$3.10 with huggingface_hub, pandas, matplotlib
A.3.4. Models
Pre-converted GGUF models for vlut.cpp are hosted at https://huggingface.co/collections/XXXXyu/vlutcpp:
For manual model conversion or baseline evaluation, use the original HuggingFace weights (see Evaluation.md).
A.4. Installation
A.4.1. Build from source.
Clone the repository and build with CMake:
git clone https://github.com/OpenBitSys/vlut.cpp.git
cd vlut.cpp
cmake -B build
cmake --build build --config Release -j4
Optional CMake flags: -DVLUT_SVE=ON for ARM SVE (e.g., AWS Graviton 3); -DTABLE_ENTRY_SIZE=<N> to change the $N$-tile size (default: 32).
A.4.2. Set up Python and download models.
Install Python dependencies and download pre-converted GGUF models. Store each model in a directory named after the model’s short name:
pip install huggingface_hub pandas matplotlib
# download pre-converted models (example: Llama3 8B)
hf download XXXXyu/Llama3-8B-1.58-100B-tokens-vlut-gguf \
ggml-model-I1_V_2.gguf ggml-model-I2_V_4.gguf \
--local-dir ~/models/Llama3-8B-1.58-100B-tokens
Repeat for other models, replacing the repository ID and -local-dir accordingly.
A.5. Experiment Workflow
A.5.1. Quick functional check.
Verify the build by running a single batched-decoding pass:
./build/bin/llama-batched \
-m ~/models/Llama3-8B-1.58-100B-tokens/ggml-model-I1_V_2.gguf \
-p "I believe" -np 32 -n 16 -t 4 -ngl 0 \
--temp 0.5 --repeat-penalty 1.5
The model should load without errors and print generated text from 32 parallel sequences.
A.5.2. Minimal reproduction (one device, one model).
Set environment variables and run the three benchmarks:
export DEVICE_NAME=mydevice
export MODEL_DIR=~/models/Llama3-8B-1.58-100B-tokens
# 1. GeMM kernel benchmark (single- and multi-threaded)
./evaluation/scripts/bench-gemm.sh "\$DEVICE_NAME" 1 256
./evaluation/scripts/bench-gemm.sh "\$DEVICE_NAME" 4 256
# 2. End-to-end prefilling
DEVICE_NAME="\$DEVICE_NAME" MODEL_DIR="\$MODEL_DIR" \
./evaluation/scripts/bench-e2e-prefill.sh
# 3. End-to-end batched decoding
DEVICE_NAME="\$DEVICE_NAME" MODEL_DIR="\$MODEL_DIR" \
./evaluation/scripts/bench-e2e-batch.sh
Raw results (CSV/log) are written to directories under evaluation/, named results_gemm_\${DEVICE_NAME}, results_e2e_prefill_\${DEVICE_NAME}, and results_e2e_batch_\${DEVICE_NAME}. The scripts benchmark all vlut.cpp quantization variants found in MODEL_DIR and skip missing baselines with warnings.
A.5.3. Other models and devices.
To evaluate additional models, change MODEL_DIR and re-run Steps 2–3 (the GeMM benchmark uses built-in shapes and only needs to run once per device). To evaluate additional devices, run on different hardware with a new DEVICE_NAME (results saved with different folder names).
A.5.4. Full paper comparison.
The same scripts automatically benchmark baselines (llama.cpp, T-MAC, bitnet.cpp) when their repositories and model files are present. Place all repositories as siblings in the same workspace and prepare baseline model files as described in Evaluation.md. Note that T-MAC and bitnet.cpp may require re-compilation per model or quantization variant.
A.6. Evaluation and Expected Results
Generate figures and summary reports from raw results. Each plotting script accepts --single-thread, --multi-thread, or --both (default):
python evaluation/scripts/plot/plot_gemm_combined.py
python evaluation/scripts/plot/plot_e2e_prefill_combined.py
python evaluation/scripts/plot/plot_e2e_batch_combined.py
Output PDF figures are saved to evaluation/figures/ and CSV reports to evaluation/reports_*/. The device identifier mydevice is recognized by the plotting scripts as a built-in dummy identifier. The paper’s canonical identifiers are: pc_intel, laptop_amd, orangepi, smartphone, aws_arm.
Absolute throughput varies by hardware, but relative speedup trends should match the paper: vlut.cpp generally outperforms baselines in GeMM throughput, prefilling speed, and parallel decoding throughput; the I1 quantization remains competitive while using fewer bits. On AWS Graviton 3, llama.cpp’s Q4_0 for HF BitNet 3B can be relatively strong due to ARM-specific 4-bit kernel optimizations. Without baselines, CSV speedup reports will be incomplete—verify raw results and figures instead.
A.7. Experiment Customization
A.7.1. Build-time options.
vlut.cpp exposes two CMake flags that affect performance:
  • -DTABLE_ENTRY_SIZE=<N>: $N$-tile size (default: 32). Larger values (e.g., 64) may improve throughput on some CPUs.
  • -DVLUT_SVE=ON: enable ARM SVE intrinsics (e.g., AWS Graviton 3).
A.7.2. Quantization types.
Five Vec-LUT packings are available (I1_V, I1_V_2, I2_V, I2_V_4, I2_V_8), controlling the $K$-tiling and packed weight layout. All five are provided as pre-converted models on HuggingFace. The examples above use I1_V_2 and I2_V_4, which are performant on most devices. To convert models with a specific packing manually, see Evaluation.md.
A.7.3. Benchmark script arguments.
The GeMM script takes positional arguments: bench-gemm.sh <device> <threads> <seq_len> [entry_size]. The end-to-end scripts accept environment variables to override defaults:
  • Prefilling (bench-e2e-prefill.sh): PROMPT_LENGTH (default 128,256,512), THREAD_COUNT (default 1,4), REPEAT_COUNT (default 3).
  • Batched decoding (bench-e2e-batch.sh): PARALLEL_SEQS (default 64,128,256), THREAD_COUNT (default 4), PREFILL_LEN (default 16), TOKEN_GEN_LENS (default 16).
A.7.4. Automated configuration search.
search-config.sh sweeps over $N$-tile sizes (16, 32, 64) and thread counts (1, 4, 8) to find the best configuration for a given device. Results are written to evaluation/results_search/.
A.7.5. Other models.
Beyond the three models evaluated in the paper, vlut.cpp also supports more ternary models. Convert and evaluate them following the instructions in Evaluation.md.
A.8. Notes
  • Set the CPU governor to performance mode and minimize background processes for stable results.
  • On heterogeneous CPUs (e.g., ARM big.LITTLE), use --cpu-mask and --cpu-strict to pin inference to performance cores.
  • The end-to-end scripts delete target result directories before writing; back up results if needed.
  • If you encounter an awk error in bench-gemm.sh, replace test-to-csv.sh with test-to-csv-backup.sh in the script.